If interest rate is unknown, give the formula for annual interest rate, j.

1
Expert's answer
2021-11-23T17:23:03-0500

We have two cases below

  1. Simple interest

"A=P(1+jn)"

where "A" is the accumulated amount, "P" is the principal amount, "j" is interest rate and "n" is the time period.

We can make "j" subject of the formula to get the unknown interest rate "j" as follows

"A=P(1+jn)"


"A=P+Pjn"


"P+Pjn=A"


"Pjn=A-P"


"{Pjn\over Pn}={A-P\over Pn}"


"\therefore j={A-P\over Pn}"


2.Compound interest

"A=P(1+j)^n"

"P(1+j)^n=A"

"{P(1+j)^n\over P}={A\over P}"

"(1+j)^n={A\over P}"

"(n\sqrt{1+j})^n=n\sqrt{{A\over P}}"

"1+j=n\sqrt{{A\over P}}"

"\therefore j=n\sqrt{{A\over P}}-1"